OpenStudy (eric_d):

Bromine consists two isotopes Br-79 n Br-81 wth relative abundance 50% n 50% respectively.Apart frm peaks at79 n 91, due to Br+ ions frm these two isotopes, the mass spectrum of r also shows peak at 158,160 n 162 a)what are the ions that give rise to these three peaks in spectrum of Br @aaronq

OpenStudy (aaronq):

they're molecular ions, 79 +79 =158, 79+81=160, 81+81=162

OpenStudy (eric_d):

Okay..what must I do next

OpenStudy (aaronq):

so just specify which isotopes make up the ions

OpenStudy (eric_d):

Br+ ?

OpenStudy (aaronq):

you're not specifying which isotope, and there are 2 (hence molecular). Basically all the combinations are there

OpenStudy (somy):

so u mean that 158 is 2 of Br 79 160 is 1 of Br 79 and 1 of Br 81 162 is 2 of Br 81

OpenStudy (somy):

like this? @aaronq

OpenStudy (aaronq):

yep, thats exactly it. He might need to put it into whatever type of notation he's using though, like \(^{79}Br_2^+\)

OpenStudy (somy):

this is how you do it @eric_d

OpenStudy (somy):

btw that 2 from right side of Br it means there are 2 of it? @aaronq

OpenStudy (aaronq):

yep, two \(^{79}Br\) isotopes

OpenStudy (eric_d):

Okay, understood Thank you

OpenStudy (aaronq):

no problem dude

OpenStudy (somy):

:D good :)